# Invertible Consistency Distillation for Text-Guided Image Editing in Around 7 Steps

### v2

@@ Abstract
11footnotemark: 1
Yandex Research                         22footnotemark: 2HSE University
Figure 1: Invertible Consistency Distillation (iCD) enables both accurate image inversion and strong generation performance in a few model evaluations.
@@ 1 Introduction
To speed-up the inference, many recent works aim to reduce the number of diffusion steps via diffusion distillation [10, 11, 12, 13, 14, 15, 16] that has provided significant progress in high-quality generation in $1{-}4$ steps and has already been successfully scaled to the state-of-the-art text-to-image diffusion models [17, 18, 19, 20, 21, 22, 23]. Though the existing distillation approaches still often trade either mode coverage or image quality for few-step inference, the proposed models can already be feasible for practical applications, such as text-driven image editing [24, 25, 26].
@@ 3.1 Forward Consistency Distillation
The transition from CD to the forward counterpart is quite straightforward: the only thing that should be modified is the boundary condition. Precisely, the forward consistency model is constrained to be an identity function for the last trajectory point: $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{N}},t_{N})=\boldsymbol{x}_{t_{N}}$. Thus, fCD inherits the same consistency distillation loss (2) without incurring extra training costs. This way, the distilled model can transform any trajectory point to the last one: $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{n}},t_{n})=\boldsymbol{x}_{t_{N}},\forall\ t_{n}$. To perform inversion, first, fCD encodes an image into noise and then the CD decodes its back. The comparison between CD and fCD is shown in Figure 3a.
@@ 3.2 Multi-boundary Consistency Distillation
Recent approaches [52, 13] generalize the CD framework to a multistep regime and allow approximating arbitrary trajectory points in the reverse direction. However, these methods focus solely on the generation quality, without supporting the inversion. Thus, inspired by these works, we propose a multi-boundary CD, that unlocks the multistep deterministic inversion with the distilled models and carries similar training costs as the classical CD methods.
Specifically, we divide the solution interval, $\{t_{0},...,t_{N}\}$, into $m$ segments and perform the distillation on each of these segments separately. This way, we obtain a set of single-step consistency models operating on different intervals and boundary points. This formulation is valid for both CD and fCD and allows for the deterministic multistep inversion. We provide an illustration of $2$-boundary CD and fCD in Figure 3b. We denote the multi-boundary reverse and forward models as CDm and fCDm.
@@
where $s^{m}_{t}$ is the boundary time step depending on the number of boundaries, $m$, and the current time step, $t$. For instance, let $m=1$, then $s^{1}_{t}=t_{0}$ for CD1 and $s^{1}_{t}=t_{N}$ for fCD1. Note that we learn a single model, the multistep is achieved by not learning parameter $s^{m}_{t}$. The training objective remains the same as (2), avoiding additional training costs compared to CD. The only limitation is that the number of segments and the corresponding boundary time steps must be set before the training.
@@ 3.3 Training fCDm and CDm
We train fCDm and CDm separately from each other but initialize them with the same teacher model. We use the same loss with a difference only in boundary time steps. However, a notable difference is the CFG scale, $w$. For CDm, we preliminarily embed the model on guidance, following [19], to use various $w$ at sampling and avoid extra model evaluations. For fCDm, we consider the unguided model with a constant $w=1$. The reason is that the guided encoding ($w>1$) leads to the out-of-distribution latent noise [31], which results in incorrect reconstruction. We confirm this intuition in Section 3.4. Finally, we find that $m{=}3{-}4$ provides good generation and inversion quality for large-scale text-to-image models [4, 1].
@@ 4.2 Text-guided image editing
In this section, we apply the proposed iCD to the text-guided image editing problem. For the SD1.5 model, we use the Prompt-to-Prompt (P2P) method [60]. We vary two hyperparameters: the cross-attention and self-attention steps balancing between editing strength and preservation of the reference image. For the SDXL model, we follow the ReNoise [24] evaluation setting and just change the source prompt during decoding according to [61].
@@ Appendix C Image inversion
In Table 5, we provide comparisons between the $3$- and $4$-step configurations of iCD-SD1.5, which perform similarly. Figure 14 shows the image inversions provided by our approach, NTI and NPI. Figure 15 shows the image inversions compared to the ReNoise method.
@@ Appendix D Text-guided image editing
Appendix E Limitations
The iCD limitations include the requirement to predetermine boundary time steps before distillation, which may restrict the model flexibility of the model. Additionally, the extra preservation losses necessitate increased computational resources during training. Furthermore, the editing method is susceptible to hyperparameter values, leading to inconsistent performance across different prompts. Finally, the quality of the current distillation techniques itself requires significant improvement, as it does not consistently achieve high standards across various scenarios.
Appendix F Broader impacts

### v3

@@ Abstract
1Yandex Research                         2HSE University
Figure 1: Invertible Consistency Distillation (iCD) enables both fast image editing and strong generation performance in a few model evaluations.
@@ 1 Introduction
To speed-up the inference, many recent works aim to reduce the number of diffusion steps via diffusion distillation [10, 11, 12, 13, 14, 15, 16, 17] that has provided significant progress in high-quality generation in $1{-}4$ steps and has already been successfully scaled to the state-of-the-art text-to-image diffusion models [18, 19, 20, 21, 22, 23, 24]. Though the existing distillation approaches still often trade either mode coverage or image quality for few-step inference, the proposed models can already be feasible for practical applications, such as text-driven image editing [25, 26, 27].
@@ 3.1 Forward Consistency Distillation
The transition from CD to the forward counterpart is quite straightforward: the only thing that should be modified is the boundary condition. Precisely, the forward consistency model is constrained to be an identity function for the last trajectory point: $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{N}},t_{N})=\boldsymbol{x}_{t_{N}}$. Thus, fCD inherits the same consistency distillation loss (2) without incurring extra training costs. This way, the distilled model can transform any trajectory point to the last one: $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{n}},t_{n})=\boldsymbol{x}_{t_{N}},\forall\ t_{n}$. To perform inversion, first, fCD encodes an image into noise and then CD decodes it back. The comparison between CD and fCD is shown in Figure 3a.
@@ 3.2 Multi-boundary Consistency Distillation
Recent approaches [53, 13] generalize the CD framework to a multistep regime and allow approximating arbitrary trajectory intervals in the reverse direction. However, these methods focus solely on the generation quality, without supporting the inversion. Thus, inspired by these works, we propose a multi-boundary CD, that unlocks deterministic multistep inversion with the distilled models and carries similar training costs as the classical CD methods.
Specifically, we divide the solution interval, $\{t_{0},...,t_{N}\}$, into $m$ segments and perform the distillation on each of these segments separately. This way, we obtain a set of single-step consistency models operating on different intervals and boundary points. This formulation is valid for both CD and fCD and hence can enable deterministic multistep inversion. We provide an illustration of $2$-boundary CD and fCD in Figure 3b. We denote the multi-boundary reverse and forward models as CDm and fCDm.
@@
where $s^{m}_{t}$ is the boundary time step depending on the number of boundaries, $m$, and the current time step $t$. For instance, let $m=1$, then $s^{1}_{t}=t_{0}$ for CD1 and $s^{1}_{t}=t_{N}$ for fCD1. Note that we learn a single model, the multistep sampling is achieved by varying $s^{m}_{t}$ during inference. The training objective remains the same as (2), avoiding additional training costs compared to CD. The only limitation is that the number of segments and the corresponding boundary time steps must be set before the training.
@@ 3.3 Training fCDm and CDm
We train fCDm and CDm separately, initializing both with the same teacher model. We use the same loss with a difference only in boundary time steps. However, a notable difference is the CFG scale, $w$. For CDm, we preliminarily embed the model on guidance, following [20], to use various $w$ during sampling and avoid extra model evaluations. For fCDm, we consider an unguided model with a constant $w=1$. The reason is that the guided encoding ($w>1$) leads to out-of-distribution latent noise [32], and as a result, to poor image reconstruction. We confirm this intuition in Section 3.4. Finally, we find that $m{=}3{-}4$ provides competitive generation and inversion quality for large-scale text-to-image models [4, 1].
@@ 4.2 Text-guided image editing
In this section, we apply the proposed iCD to the text-guided image editing problem. For the SD1.5 model, we use the Prompt-to-Prompt (P2P) method [61]. We vary two hyperparameters: the cross-attention and self-attention steps balancing between editing strength and preservation of the reference image. We also apply our approach to MasaCTRL [62] in Appendix D. For the SDXL model, we follow the ReNoise [25] evaluation setting and just change the source prompt during decoding according to [63].
@@ Appendix B Image generation with iCD
Figure 13: Comparison between Prompt2Prompt and MasaCTRL applied to our approach
@@ Appendix C Image inversion
Method Ours 8 steps, SD1.5 NTI, SD1.5 NPI, SD1.5 Ours 8 steps, SDXL ReNoise, LCM-XL
Time, secs $0.959\pm.005$ $116.4\pm.1$ $9.95\pm.03$ $1.56\pm.07$ $6.75\pm.52$
Table 6: The time required to invert a single image for different approaches.
@@
In Table 5, we provide comparisons between the $3$- and $4$-step configurations of iCD-SD1.5, which perform similarly. Figure 14 shows the image inversions provided by our approach, NTI and NPI. Figure 15 shows the image inversions compared to the ReNoise method. In Table 6, we present the time required to invert a single image for different methods.
@@ Appendix D Text-guided image editing
Figure 13 provides a few examples of our approach combined with MasaCTRL [62] for non-rigid editing. The results demonstrate that our method can be used with various editing methods.
We also present the comparison with SDXL-Turbo using SDEdit [63] in Figure 20 (upper). We observe that SDXL Turbo significantly hurts reference image preservation due to stochasticity. This highlights the importance of accurate image inversion for editing. Moreover, we compare our approach with Intruct-Pix2Pix [7], Figure 20 (bottom), and observe that it outperforms Intruct-Pix2Pix in terms of both content preservation and editing strength while being training-free.
@@
Figure 20: Image editing results produced by our method and SDXL-turbo using SDEdit (upper) and Instruct-P2P (bottom).
Appendix E Failure cases
We present some inversion failure cases in Figure 21. Our method sometimes oversaturates images for high guidance scales and struggles to reconstruct complex details like human faces and hands
Figure 21: Failure cases.
Appendix F Limitations
The iCD limitations include the requirement to predetermine boundary time steps before distillation, which may restrict the model flexibility. Additionally, the extra preservation losses necessitate increased computational resources during training. Furthermore, the editing method is susceptible to hyperparameter values, leading to inconsistent performance across different prompts. Finally, the quality of the current distillation techniques itself requires significant improvement, as it does not consistently achieve high standards across various scenarios.
Appendix G Broader impacts